estimate the average drift speed of conduction electron in a copper wire of crossection area 1*carrying a current of 1.5A .Assume that each cu atom contributes roughly one conduction electron .the density of cu is and its atomic mass is 63.5u ans--1.1mms^-1

explain with step wise solution

thanks


Grade:upto college level

1 Answers

Utkarsh Singh
5 years ago
Home >> JEEMAIN and NEET >> Physics >> Class12 >> Current ElectricityWhat is the average drift speed of conduction electrons in a copper wire of cross-sectional area 1.0×10–7m2 carrying a current of 1.5A. Assume that each copper atom contributes roughly one conduction electron. The density of copper is 9.0×103kg/m3 , and its atomic mass is 63.5u.(A)1.9mm/s(C)2.9mm/s(b)1.1mm/s(D)2.2mm/s jeemain physics current-electricity cbse class12 ch5 electric-current-concepts difficult Clay6 Question Share Shareasked Feb 14, 2014 by thanvigandhi_1 edited Sep 4, 2014 by thagee.vedartham1 AnswerNeed homework help? Click here.The direction of drift velocity of conduction electrons is opposite to the electric field direction,i.e., electrons drift in the direction of increasing potential. The drift speedvd is given byvd=(I/neA)Now e=1.6×10−19C,A=1.0×10−7m2,I=1.5AThe density of conduction electrons, n is equal to the number of atoms per cubic meter(assuming one conduction electron per Cu atom as is reasonable from its valence electron count of one).A cubic meter of copper has a mass of 9.0×103kg. Since 6.0×1023 copper atoms have a mass of 63.5g,n=6×102363.5 ×9×106=8.5×1028on substituting in the formula forvd, one gets the final answer as 1.1mm/s